OpenStudy (anonymous):

true or false the solution of x+5/2=15 is 25

OpenStudy (anonymous):

x=15 -5/2 x=\[\frac{ 30-5 }{ 2 }=\frac{ 25 }{ 2 }\] so FALSE

OpenStudy (raffle_snaffle):

it's true

OpenStudy (raffle_snaffle):

oh shiz sorry you are right...

OpenStudy (raffle_snaffle):

no it's true

OpenStudy (raffle_snaffle):

i just solved it by hand and used my calculator

OpenStudy (raffle_snaffle):